# Axion Dark Matter Archaeology with Primordial Gravitational Waves

### v1

@@ 2.1 Condition for heavy quark domination
$\langle\sigma v\rangle_{\bar{Q}Q}=\frac{\pi\alpha_{s}^{2}}{16m_{Q}^{2}}\left(c_{f}n_{f}+c_{g}\right).$ (2.5)
[…]
$T_{\rm dom}\approx 10^{7}\left(\frac{m_{Q}}{10^{12}\,\text{GeV}}\right)^{2}\,\text{GeV}.$ (2.6)
In order for heavy quarks to dominate the early Universe, their decay must occur after heavy quark-radiation equality. Since the plasma temperature is always decreasing, we require $T_{\rm decay}<T_{\rm dom}$. To estimate $T^{Q}_{\rm decay}$, we set $\Gamma_{Q}\equiv H_{\rm rad}\left(T=T^{Q}_{\rm decay}\right)$ where $H_{\rm rad}$ is the Hubble parameter in the radiation-dominated Universe,
$T_{\rm decay}^{d}=\left(\Gamma^{d}_{Q}\right)^{\frac{1}{2}}\left(\frac{1}{g_{\star}(T_{\rm decay})}\right)^{\frac{1}{4}}\left(\frac{90\,m_{\rm Pl}^{2}}{8\pi^{3}}\right)^{\frac{1}{4}}.$ (2.7)
@@
$\Gamma_{Q}^{d=6}$ $=\frac{m_{Q}^{5}}{512\pi^{3}\Lambda^{4}},\,\,\,\,$ $\Gamma_{Q}^{d=7}$ $=\frac{m_{Q}^{7}}{49152\pi^{5}\Lambda^{6}},\,$ (2.8)
where we neglect the details of higher dimension decays in this paper. Furthermore, the main focus in this article will be the dimension 6 and 7 dimensional decay terms. This is because $d\leq 5$ models will not produce periods of HQD that will alter the misalignment mechanism. Models with $d=5$ decays can lead to HQD, but these typically require $m_{Q}>f_{a}$ or predict overproduction of the axion due to a large $f_{a}$ and has no in-built mechanism to dilute it. Using Eq. (2.8) one can estimate when $T_{\rm dom}>T_{\rm decay}$. For $d=6$ and $d=7$, this is when
$\Lambda\gtrsim(4\times 10^{28}\,{\rm GeV}^{2}\,\times m_{Q}M_{\rm Pl})^{1/4}\,\,\,\,{\rm and}\,\,\,\,\Lambda\gtrsim(4\times 10^{25}\,{\rm GeV}^{2}\,\times m_{Q}^{3}M_{\rm Pl})^{1/6},$ (2.9)
respectively. Furthermore, with the calculation of $T_{\rm dec}$ we are able to assess whether a particular model and parameter point is in tension with BBN [84, 85, 86, 87]. Typically one estimates this by checking that $T_{\rm dec}\gtrsim 3\,{\rm MeV}$.
2.2 Axion misalignment with heavy quark domination
The details of the axion misalignment mechanism within non-standard cosmologies can be found in Refs. [74, 75, 76, 77, 78, 79, 80, 81, 88]. The insight of Ref. [40] was that axion models themselves typically contain all necessary ingredients for early matter domination.
@@ 2.2 Axion misalignment with heavy quark domination
where $\rho$, $n$ and $s$ are the energy, number, and entropy densities respectively, we also denote each component, SM radiation (R), heavy quark ($Q$), in the subscript. Note that we haven’t included the Boltzmann equation for the the thermally produced axion. As shown in Refs. [40, 89], if HQD occurs, the branching ratio to axions must be subdominant otherwise axion contributions to $\Delta N_{\rm eff}$ would be too large. Hence, in this work we only consider 100% decay into SM particles, which implies that if HQD occurs and ends after thermal axion decoupling, then $\Delta N_{\rm eff}$ will be much less than the expected value from the standard thermal axion, i.e. $\Delta N_{\rm eff}\ll 0.027$. This is despite the axion actually having been in thermal contact with the SM early on.
One can solve Eqs. (2.10) to validate the approximations made in Eq. (2.6) and Eq. (2.7), and to when generate the GW signals in section 3. We use the precise determination of $H$ as an input to solve the misalignment mechanism using the dedicated solver MiMeS [90]. Specifically, we solve the equation of motion for the axion field, $a$, which is expressed in terms of the axion angle, $\theta=a/f_{a}$
@@
where $\tilde{m}_{a}(t)$ is the thermal mass of the axion [91]. Generically, the axion abundance is diluted such that a greater $f_{a}$ is required to achieve the correct relic abundance for axion dark matter.
@@
In order to make the connection with the experimental search for axion dark matter, we numerically solve Eqs. (2.10) and Eq. (2.11) to find the required $f_{a}$ such that $\Omega_{a}\,h^{2}=\Omega_{\rm DM}\,h^{2}$, where we take the relic abundance of dark matter determined in [92]. We do this for a range of $m_{Q}$ and $\Lambda$ values, and take $\theta_{i}=2.15$ in-line with calculations of post-inflationary PQ breaking [73]. Furthermore, we check that $\rho_{\rm R}>\rho_{Q}$ at the onset of BBN. In previous works $m_{Q}\sim f_{a}$ and $\Lambda\sim m_{\rm Pl}$ were taken, but these are simply the upper bounds on such parameters, we find that by insisting that the QCD axion has to account for all of the dark matter, these assumptions can be relaxed and we still have a sufficiently constrained scenario.
@@
Figure 1: The range of axion mass $m_{a}$ which can reproduce the correct relic abundance allowing for any $m_{Q}\lesssim f_{a}$ and $\Lambda<m_{\rm Pl}$. This is for $d=6$ but we find the result for $d=7$ decays to be very similar. The axion limits and QCD axion band was plotted using Ref. [93].
@@
Fig. 1 shows the range of $m_{a}$ values that can achieve the correct relic abundance, allowing $m_{Q}$ and $\Lambda$ to vary. We have presented the possible $m_{a}$ range in the $\left(m_{a},g_{a\gamma}\right)$-plane, where $g_{a\gamma}$ is the axion-photon coupling which depends on the anomaly coefficients, $N$, from Eq. (2.1) and $E$, the coefficient of the electromagnetic anomaly term,
$g_{a\gamma}\equiv\frac{\alpha}{2\pi}\frac{1}{f_{a}}\left(\frac{E}{N}-1.92(4)\right).$ (2.12)
@@
$\alpha$ is the electromagnetic fine-structure constant. In this article, we remain agnostic about the specific axion model and therefore show a representative line for $E/N=0$. Furthermore, the yellow band shows a model independent band encompassing most QCD axion models. We note that by allowing $m_{Q}\lesssim f_{a}$ we have found a substantial quantitative difference from the results in Refs. [40, 41] where models with $d=6$ decay terms were found to only marginally change the expected $m_{a}$ value that reproduces the correct relic density, even when $\Lambda$ could vary. In Fig. 1 we see that $d=6$ decays can alter the cosmology such that the $f_{a}$ required for $\Omega_{a}\,h^{2}=\Omega_{\rm DM}\,h^{2}$ is orders of magnitude higher. What limits this value is the constraints from BBN. The left most point on the axion dark matter line corresponds to $f_{a}=4\times 10^{14}\,{\rm GeV}$, $m_{Q}=4\times 10^{11}\,{\rm GeV}$, and $\Lambda=4\times 10^{18}\,{\rm GeV}$. Calculating $T_{\rm dec}$ via Eq. (2.7) gives $5\,{\rm MeV}$, but due the more precise numerical evaluation from Eq. (2.10) gives $T_{\rm dec}=3\,{\rm MeV}$.
@@
We show this slight difference in the way we calculate the constraint from BBN on the $\left(m_{Q},\Lambda\right)$-plane in Fig. 2. The black and white hatched region is bounded by a dotted and a solid line, which indicates the approximate bound (Eq. (2.7)) or full numerical one (Eqs. (2.10)) respectively. One can see the constraints coming from the more accurate numerical evaluation is more aggressive, henceforth the BBN bounds we will show correspond to this. All the points in Fig. 2 to the right of the hatched region can satisfy $\Omega_{a}=\Omega_{\rm DM}$ given the certain choice of $f_{a}$. Some of those $f_{a}$ values will coincide with the haloscope experiment constraints shown in Fig. 2 as a dark gray region. In particular, the two experiments that have reached sensitivities down to the KSVZ line in the relevant $f_{a}$ region is ADMX [16, 94] and CAPP [95, 96, 18, 97], they are the crimson red and orange in shaded regions in Fig. 2. Note that since these experiments are haloscope experiments, their sensitivity is intimately tied to the $\Omega_{a}=\Omega_{\rm DM}$ requirement we made. In the left panel of Fig. 2 we show the results for $d=6$ decay terms, and in the right panel we show the results for $d=7$ decay terms. This will be useful in assessing the potential for future GW experiments to probe the same parameter space. As we will show in section 3, the inflationary GW signal is sensitive $m_{Q}$ and $\Lambda$ and not $f_{a}$ directly.
@@
Figure 2: The current constraints on the $m_{Q}$ and $\Lambda$ parameter space from ADMX and CAPP experiments. The left panel is for $d=6$ decays and the right panel is for $d=7$ decays. The value of $f_{a}$ is not a priori set, instead we determine the required value such that $\Omega_{a}=\Omega_{\rm cdm}$. The hatched region is where $Q$ decay occurs too late and will be constrained by BBN, we have bounded it by the dotted and solid black lines which correspond to the more approximate Eq. (2.7) and precise numical evaluation via Eqs. (2.10) respectively.
@@
Before we move on, we comment that we have not included any contribution to the axion abundance from topological defects such as cosmic strings or unstable domain walls [98, 99, 100]. We do this because the exact contribution is still under dispute in the literature and all calculations to date have been performed under standard cosmology [101, 102, 103, 104, 105, 106, 107]. Furthermore, since the vast majority of the axion abundance is produced around the QCD phase transition $T_{\rm QCD}$, similarly to the misalignment mechanism, the effect where HQD dilutes the axion abundance will also hold. Future work is warranted in this direction, but it is beyond the scope of this article.
2.3 Interpreting the cosmology of a low mass axion
@@
In the previous subsection we have shown that axion models with post-inflationary PQ breaking can lead to axion dark matter with masses as low as $\sim 10^{-8}\,{\rm eV}$. The only requirements are that the new particles that generate the anomaly term in Eq. (2.2), are heavy ($m_{Q}\gtrsim 10^{7}\,{\rm GeV}$) and live long enough. For the effective parametrisation of Eq. (2.4), this translates to the requirement that the lowest dimension decay terms have to be $d\geq 6$. What’s more, recently Ref. [41] found that there exists four dimension 6 KSVZ-type charge configurations that do not have a domain wall problem, for lower dimensions, there are only two configurations. Additionally, it appears that these new models contain $Q$ decays to SM particles only, therefore $\Delta N_{\rm eff}\ll 0.027$ [89].
@@
As always, we would like to find experimental avenues to break the degeneracy between the pre and post-inflationary breaking scenario. What’s more, some kind of understanding of when $U(1)_{\rm PQ}$ breaking occurred relative to the reheating temperature would constitute a profound insight into the nature and scale of inflation. As we’ve seen in Eq. (2.6) and Eq. (2.8) the onset and end of HQD is sensitive to fundamental parameters of the underlying theory, so if we can probe these events experimentally in some way, we will be able to learn about the possible axion models. As we will see in the next section, inflationary gravitational waves may provide precisely this insight.
3 Primordial Gravitational Waves

### v2

@@ II.1 Condition for heavy quark domination
$\langle\sigma v\rangle_{\bar{Q}Q}=\frac{\pi\alpha_{s}^{2}}{16m_{Q}^{2}}\left(c_{f}n_{f}+c_{g}\right).$ (6)
where $n_{f}$ is the number of quark flavours that $Q$ can annihilate into, and $\left(c_{f},c_{g}\right)=\left(2/9,\,220/27\right)$ for triplets and $\left(c_{f},c_{g}\right)=\left(2/3,\,27/4\right)$ for octets, in this paper we assume the former. Through the standard thermal freeze-out procedure one can determine the temperature of the SM plasma when $\rho_{Q}\approx\rho_{\rm SM}$ Cheek et al. (2024),
$T_{\rm dom}\approx 10^{7}\left(\frac{m_{Q}}{10^{12}\,\text{GeV}}\right)^{2}\,\text{GeV}.$ (7)
In order for heavy quarks to dominate the early Universe, their decay must occur after heavy quark-radiation equality. Since the plasma temperature is always decreasing, we require $T_{\rm dec}<T_{\rm dom}$. To estimate $T^{Q}_{\rm dec}$, we set $\Gamma_{Q}\equiv H_{\rm rad}\left(T=T^{Q}_{\rm dec}\right)$ where $H_{\rm rad}$ is the Hubble parameter in the radiation-dominated Universe [^1] Notice that we have assumed $H_{\rm rad}$ for this approximation, which strictly speaking, is not appropriate during early matter domination. However, when the matter domination is caused by the decaying particle, this estimate holds up-to $\mathcal{O}(1)$ factors. In the later analysis of this article we solve the coupled Boltzmann equations, which show broad consistency with this estimate.,
$T_{\rm dec}^{d}\approx\left(\Gamma^{d}_{Q}\right)^{\frac{1}{2}}\left(\frac{1}{g_{\star}(T_{\rm dec})}\right)^{\frac{1}{4}}\left(\frac{90\,m_{\rm Pl}^{2}}{8\pi^{3}}\right)^{\frac{1}{4}}.$ (8)
@@
$\Gamma_{Q}^{d=6}$ $=\frac{m_{Q}^{5}}{512\pi^{3}\Lambda^{4}},\,\,\,\,$ $\Gamma_{Q}^{d=7}$ $=\frac{m_{Q}^{7}}{49152\pi^{5}\Lambda^{6}},\,$ (9)
where we neglect the details of higher dimension decays in this paper. Furthermore, the main focus in this article will be the dimension 6 and 7 dimensional decay terms. This is because $d\leq 5$ models will not produce periods of HQD that will alter the misalignment mechanism. Models with $d=5$ decays can lead to HQD, but these typically require $m_{Q}>f_{a}$ or predict overproduction of the axion due to a large $f_{a}$ and has no in-built mechanism to dilute it. According to Ref. Di Luzio et al. (2024), there are 43 models that have dimension 6 decays at lowest order and 44 for dimension 7. Using Eq. (9) one can estimate when $T_{\rm dom}>T_{\rm dec}$. For $d=6$ and $d=7$, this is when
$\Lambda$ $\gtrsim$ $(4\times 10^{28}\,{\rm GeV}^{2}\,\times m_{Q}M_{\rm Pl})^{1/4}\,\,\,\,{\rm and}$
$\Lambda$ $\gtrsim$ $(4\times 10^{25}\,{\rm GeV}^{2}\,\times m_{Q}^{3}M_{\rm Pl})^{1/6},$ (10)
respectively. Furthermore, with the calculation of $T_{\rm dec}$ we are able to assess whether a particular model and parameter point is in tension with BBN Kawasaki et al. (2005); Jedamzik (2006, 2008); Kawasaki et al. (2018). Typically one estimates this by checking that $T_{\rm dec}\gtrsim 3\,{\rm MeV}$.
II.2 Axion misalignment with heavy quark domination
The details of the axion misalignment mechanism within non-standard cosmologies can be found in Refs. Steinhardt and Turner (1983); Lazarides et al. (1990); Kawasaki et al. (1996); Visinelli and Gondolo (2010); Nelson and Xiao (2018); Ramberg and Visinelli (2019); Arias et al. (2021); Xu (2023); Chen et al. (2025). The insight of Ref. Cheek et al. (2024) was that axion models themselves typically contain all necessary ingredients for early matter domination.
@@ II.2 Axion misalignment with heavy quark domination
where $\rho$, $n$ and $s$ are the energy, number, and entropy densities respectively, we also denote each component, SM radiation (R), heavy quark ($Q$), in the subscript. Note that we haven’t included the Boltzmann equation for the the thermally produced axion. As shown in Refs. Cheek et al. (2024); Cheek and Min (2025), if HQD occurs, the branching ratio to axions must be subdominant otherwise axion contributions to $\Delta N_{\rm eff}$ would be too large. Hence, in this work we only consider 100% decay into SM particles, which implies that if HQD occurs and ends after thermal axion decoupling, then $\Delta N_{\rm eff}$ will be much less than the expected value from the standard thermal axion, i.e. $\Delta N_{\rm eff}\ll 0.027$. This is despite the axion actually having been in thermal contact with the SM early on.
One can solve Eqs. (11) to validate the approximations made in Eq. (7) and Eq. (8), and to when generate the GW signals in section III. We use the precise determination of $H$ as an input to solve the misalignment mechanism using the dedicated solver MiMeS Karamitros (2022). Specifically, we solve the equation of motion for the axion field, $a$, which is expressed in terms of the axion angle, $\theta=a/f_{a}$
@@
where $\tilde{m}_{a}(t)$ is the thermal mass of the axion Borsanyi et al. (2016). Generically, the axion abundance is diluted such that a greater $f_{a}$ is required to achieve the correct relic abundance for axion dark matter.
@@
In order to make the connection with the experimental search for axion dark matter, we numerically solve Eqs. (11) and Eq. (12) to find the required $f_{a}$ such that $\Omega_{a}\,h^{2}=\Omega_{\rm DM}\,h^{2}$, where we take the relic abundance of dark matter determined in Aghanim et al. (2020). We do this for a range of $m_{Q}$ and $\Lambda$ values, and take $\theta_{i}=2.15$ in-line with calculations of post-inflationary PQ breaking Grilli di Cortona et al. (2016). Furthermore, we check that $\rho_{\rm R}>\rho_{Q}$ at the onset of BBN. In previous works $m_{Q}\sim f_{a}$ and $\Lambda\sim m_{\rm Pl}$ were taken, but these are simply the upper bounds on such parameters, we find that by insisting that the QCD axion has to account for all of the dark matter, these assumptions can be relaxed and we still have a sufficiently constrained scenario.
@@
Figure 1: The range of axion mass $m_{a}$ which can reproduce the correct relic abundance allowing for any $m_{Q}\lesssim f_{a}$ and $\Lambda<m_{\rm Pl}$. This is for $d=6$ but we find the result for $d=7$ decays to be very similar. The axion limits and QCD axion band was plotted using Ref. O’Hare (2020).
@@
Fig. 1 shows the range of $m_{a}$ values that can achieve the correct relic abundance, allowing $m_{Q}$ and $\Lambda$ to vary. We have presented the possible $m_{a}$ range in the $\left(m_{a},g_{a\gamma}\right)$-plane, where $g_{a\gamma}$ is the axion-photon coupling which depends on the anomaly coefficients, $N$, from Eq. (2) and $E$, the coefficient of the electromagnetic anomaly term,
$g_{a\gamma}\equiv\frac{\alpha}{2\pi}\frac{1}{f_{a}}\left(\frac{E}{N}-1.92(4)\right).$ (13)
@@
$\alpha$ is the electromagnetic fine-structure constant. The precise value for $E/N$ depends on the SM and PQ charges of the HQs, in other words the specific axion model one is considering. To encompass the range of $E/N$ values we show a yellow band. To represent the range of $m_{a}$ where the correct relic abundance is achieved we chose the original KSVZ model, $E/N=0$, which is the black line connecting the two black points. However, we emphasize that axion models of interest are in the range of $E/N$ values that lead to the range $\left|E/N-1.92(4)\right|\in\left[0.003-100\right]$. We note that by allowing $m_{Q}\lesssim f_{a}$ we have found a substantial quantitative difference from the results in Refs. Cheek et al. (2024); Di Luzio et al. (2024) where models with $d=6$ decay terms were found to only marginally change the expected $m_{a}$ value that reproduces the correct relic density, even when $\Lambda$ could vary. In Fig. 1 we see that $d=6$ decays can alter the cosmology such that the $f_{a}$ required for $\Omega_{a}\,h^{2}=\Omega_{\rm DM}\,h^{2}$ is orders of magnitude higher. What limits this value is the constraints from BBN. The left most point on the axion dark matter line corresponds to $f_{a}=4\times 10^{14}\,{\rm GeV}$, $m_{Q}=4\times 10^{11}\,{\rm GeV}$, and $\Lambda=4\times 10^{18}\,{\rm GeV}$. The right most point corresponds to the expectation for a dark matter axion without HQD. Calculating $T_{\rm dec}$ via Eq. (8) gives $5\,{\rm MeV}$, but due the more precise numerical evaluation from Eq. (11) gives $T_{\rm dec}=3\,{\rm MeV}$.
@@
We show this slight difference in the way we calculate the constraint from BBN on the $\left(m_{Q},\Lambda\right)$-plane in Fig. 2. The black and white hatched region is bounded by a dotted and a solid line, which indicates the approximate bound (Eq. (8)) or full numerical one (Eqs. (11)) respectively. One can see the constraints coming from the more accurate numerical evaluation is more aggressive, henceforth the BBN bounds we will show correspond to this. Furthermore, a more accurate determination is possible by combining our Boltzmann equations with those of the light elements Kawasaki et al. (2018), we leave this for future work. All the points in Fig. 2 to the right of the hatched region can satisfy $\Omega_{a}=\Omega_{\rm DM}$ given the certain choice of $f_{a}$. Some of those $f_{a}$ values will coincide with the haloscope experiment constraints shown in Fig. 2 as a dark gray region. In particular, the two experiments that have reached sensitivities down to the KSVZ line in the relevant $f_{a}$ region is ADMX Asztalos et al. (2010); Stern (2016) and CAPP Kwon et al. (2021); Yang et al. (2023); Kim et al. (2024a); Ahn et al. (2024), they are the crimson red and orange in shaded regions in Fig. 2. Note that since these experiments are haloscope experiments, their sensitivity is intimately tied to the $\Omega_{a}=\Omega_{\rm DM}$ requirement we made. In the left panel of Fig. 2 we show the results for $d=6$ decay terms, and in the right panel we show the results for $d=7$ decay terms. This will be useful in assessing the potential for future GW experiments to probe the same parameter space. As we will show in section III, the inflationary GW signal is sensitive $m_{Q}$ and $\Lambda$ and not $f_{a}$ directly.
@@
Figure 2: The current constraints on the $m_{Q}$ and $\Lambda$ parameter space from ADMX and CAPP experiments. The left panel is for $d=6$ decays and the right panel is for $d=7$ decays. The value of $f_{a}$ is not a priori set, instead we determine the required value such that $\Omega_{a}=\Omega_{\rm cdm}$. The hatched region is where $Q$ decay occurs too late and will be constrained by BBN, we have bounded it by the dotted and solid black lines which correspond to the more approximate Eq. (8) and precise numerical evaluation via Eqs. (11) respectively.
@@
Before we move on, we comment that we have not included any contribution to the axion abundance from topological defects such as cosmic strings or unstable domain walls Kibble (1976); Kibble et al. (1982); Vilenkin and Everett (1982). We do this because the exact contribution is still under dispute in the literature and all calculations to date have been performed under standard cosmology Hagmann et al. (2001); Wantz and Shellard (2010); Hiramatsu et al. (2011); Kawasaki et al. (2015); Buschmann et al. (2022); Gorghetto et al. (2021); Kim et al. (2024b). Furthermore, since the vast majority of the axion abundance is produced around the QCD phase transition $T_{\rm QCD}$, similarly to the misalignment mechanism, the effect where HQD dilutes the axion abundance will also hold. Future work is warranted in this direction, but it is beyond the scope of this article.
II.3 Interpreting the cosmology of a low mass axion
@@
In the previous subsection we have shown that axion models with post-inflationary PQ breaking can lead to axion dark matter with masses as low as $\sim 10^{-8}\,{\rm eV}$. The only requirements are that the new particles that generate the anomaly term in Eq. (3), are heavy ($m_{Q}\gtrsim 10^{7}\,{\rm GeV}$) and live long enough. For the effective parametrisation of Eq. (5), this translates to the requirement that the lowest dimension decay terms have to be $d\geq 6$. What’s more, recently Ref. Di Luzio et al. (2024) found that there exists four dimension 6 KSVZ-type charge configurations that do not have a domain wall problem, for lower dimensions, there are only two configurations. Additionally, it appears that these new models contain $Q$ decays to SM particles only, therefore $\Delta N_{\rm eff}\ll 0.027$ Cheek and Min (2025).
@@
As always, we would like to find experimental avenues to break the degeneracy between the pre and post-inflationary breaking scenario. What’s more, some kind of understanding of when $U(1)_{\rm PQ}$ breaking occurred relative to the reheating temperature would constitute a profound insight into the nature and scale of inflation. As we’ve seen in Eq. (7) and Eq. (9) the onset and end of HQD is sensitive to fundamental parameters of the underlying theory, so if we can probe these events experimentally in some way, we will be able to learn about the possible axion models. As we will see in the next section, inflationary gravitational waves may provide precisely this insight.
III Primordial Gravitational Waves